OpenStudy (fools101):

Other please help!! y = x2 + 4x + 8

OpenStudy (anonymous):

Possible intermediate steps: x^2+4 x+8 = 0 Subtract 8 from both sides: x^2+4 x = -8 Add 4 to both sides: x^2+4 x+4 = -4 Factor the left hand side: (x+2)^2 = -4 Take the square root of both sides: sqrt(x+2) = 2 i Eliminate the absolute value: x+2 = -2 i or x+2 = 2 i Subtract 2 from both sides: x = -2-2 i or x+2 = 2 i Subtract 2 from both sides: x = -2-2 i or x = -2+2 i

OpenStudy (fools101):

thank u also need help with this one please help y = x2 - 2x + 5

OpenStudy (anonymous):

Possible intermediate steps: x^2-2 x+5 = 0 Subtract 5 from both sides: x^2-2 x = -5 Add 1 to both sides: x^2-2 x+1 = -4 Factor the left hand side: (x-1)^2 = -4 Take the square root of both sides: sqrt(x-1) = 2 i Eliminate the absolute value: x-1 = -2 i or x-1 = 2 i Add 1 to both sides: x = 1-2 i or x-1 = 2 i Add 1 to both sides: x = 1-2 i or x = 1+2 i

OpenStudy (anonymous):

is this ur ans!!!!!

OpenStudy (fools101):

ohh kk thank u!!...